# LLM-Craft: Robotic Crafting of Elasto-Plastic Objects with Large Language Models

### v1

@@ 4.5 Semantic Goals
Up until this point, all of our experiments as well as past literature for elasto-plastic object robotic crafting has focused on explicit shape goals [2, 4, 1, 17, 3]. However, the core goal of creating a sculpture is to represent a shape or an idea, not not necessarily exactly replicate an existing sculpture or object. These more semantic goals are incredibly difficult to capture, quantify and evaluate, which is why to the best of our knowledge no prior work has explored this. However, given that we are using language in the LLM-Craft framework, we have a system that is more aligned with understanding and creating these semantic-based goals. For example, the termination prompt is evaluating more abstract shape similarity as opposed to low-level point similarity. We conducted a small set of experiments to evaluate if and how LLM-Craft handles more ambiguous semantic goals. We modified the action prompt to remove the inclusion of the goal image and description and instead included semantic commands such as ”make it thinner”, ”make it bumpy”, or ”make it shorter”. The qualitative results of these experiments are visualized in Figure 6. We found that the LLM-Craft framework was indeed able to modify the starting shape to better match these semantic goals. Even with the restricted action space, we are able to visually see the difference in how LLM-Craft chose to change the starting state to reach these semantically different goals.
5 Conclusion
@@
In this work, we present LLM-Craft, to the best of our knowledge the first LLM-based robotic shape crafting system. Through real-world experiments, we demonstrate the powerful reasoning capabilities of LLMs out-of-the-box for both single-step and longer horizon shape goals. However, we find that while the semantic reasoning capabilities of LLMs are very powerful, LLM-based systems sometimes hallucinate bad actions. In the LLM-Craft system there is no real-world grounding, meaning sometimes the LLM selects actions and the explanation of the predicted effect of the action does not make any physical sense. However, we believe the semantic reasoning capabilities of LLMs could be very useful in future work combining the more classical and physics-based approaches with LLMs for the higher-level reasoning and semantic similarity analysis.
Limitations. One key limitation of LLM-Craft is that the simplified state and action space limits the final shape quality of the sculptures, as even the human baseline struggled to create more detailed features. Additionally, the LLM-Craft system would sometimes make mistakes with respect to rotation between the final shape and the goal shape because the system is reasoning about overall shape similarity which may not include rotation. Finally, there currently is no true enforcement of physics and the LLM can sometimes predict inaccurate behaviors. In future work, we hope to explore incorporating LLM-based shape-level reasoning with a more traditional dynamics-based system to address these concerns.
Acknowledgments
Alison Bartsch is supported by the Philip and Marsha Dowd Engineering Seed Fund research fellowship.
Appendix A Supplementary Materials
A.1 Prompting Details
In the following sections, we show the exact prompts used for the different variations of LLM-Craft experiments, including the primary action prompt, the single-step prompts with image and array state/goal variations, the termination prompt, the semantic goal prompt, and the prompt ablations.
A.1.1 General Multi-Step Prompt
A.2 Single-Step Prompt Gridded Image Representation
A.2.1 Single-Step Prompt Binary Array Representation
A.2.2 Termination Checker Prompt
A.3 Semantic Goals Prompt Details
A.4 Prompt Ablation Details
A.5 LLM Reasoning Detailed Examples
We found that the LLM reasoned about the clay state and goal at the shape level, understanding the high-level changes that needed to be made to reach the goal. For concrete examples along with the state and goals that the LLM is referencing, please see Figure 13 above.
A.6 More Details of Action Transformation
We secure the clay to the table with a small 3D printed screw that is attached to the elevated stage. This screw resists the small upward force applied to the clay if it sticks to the fingertips while the robot is moving to the observation pose, and ensures that our assumption that the clay remains on the table is true. We know the real-world position of this 3D printed screw and set that position to be the center point of the 4x4 grid with each grid cell being 2cm x 2cm. Thus, given a discrete action, for example squeeze A1 and B2, then we know the ground truth x, y, z position of the center of both of those cells. We then set the position of the end-effector to the mid-point between the two cells and the rotation to align with the direction vector pointing from cell1 to cell2. Finally, we apply a squeeze of a final fixed distance between the fingers of 1.5 cm.
A.7 Discussion of Failure Modes
While we have demonstrated that the LLM-Craft framework can successfully create a variety of simple shapes in clay, the system can encounter a few different types of failure modes as listed below.
1.
Incorrectly predicting the clay behavior: LLM-Craft is not grounded with real-world dynamics, and thus can make mistakes predicting behavior that is not physically possible. For example, sometimes when explaining the predicted effect of a squeeze such as cells A1 and B2, the LLM will correctly predict that this will push some clay out of both cells and towards each other, but forget that some clay will also be pushed out perpendicular to the squeeze into the neighboring cells. This could be remedied in future work by incorporating a dynamics model into the framework.
2.
Entering irreversible clay states: While the LLM-Craft framework uses an iterative replanning scheme to be able to adapt crafting sequences and fix mistakes, due to the nature of clay and the action type of squeezing only, there are certain clay states that the system can create that are impossible to undo. This means that if the agent selects a bad action and it is physically impossible to reach the goal from this new clay state with only grasping-based actions, then the entire rollout will be a failure. This could be addressed in future work by incorporating additional types of actions that enable resetting the clay state.
3.
Makes mistakes with regard to the grid representation: The LLM can sometimes get confused about the grid representation and while it is clear the higher level reasoning is correct, the actual actions the LLM selects are wrong given the described target outcome. For example, the LLM agent may describe wanting to pinch the upper right corner of the clay which is the correct choice given the current state and goal, but select the cells A2 and B1, which would be pinching the upper left corner.

### v2

@@ IV-F Semantic Goals
Up until this point, all of our experiments as well as past literature for elasto-plastic object robotic crafting has focused on explicit shape goals [bartsch2023sculptbot, bartsch2024sculptdiff, shi2022robocraft, bauer2024doughnet, shi2023robocook]. However, the core goal of creating a sculpture is to represent a shape or an idea, not not necessarily exactly replicate an existing sculpture or object. These more semantic goals are incredibly difficult to capture, quantify and evaluate, which is why to the best of our knowledge no prior work has explored this. However, given that we are using language in the LLM-Craft framework, we have a system that is more aligned with understanding and creating these semantic-based goals. For example, the termination prompt is evaluating more abstract shape similarity as opposed to low-level point similarity. We conducted a small set of experiments to evaluate if and how LLM-Craft handles more ambiguous semantic goals. We modified the action prompt to remove the inclusion of the goal image and description and instead included semantic commands such as ”make it thinner”, ”make it bumpy”, or ”make it shorter”. The qualitative results of these experiments are visualized in Figure 8. We found that the LLM-Craft framework was indeed able to modify the starting shape to better match these semantic goals. Even with the restricted action space, we are able to visually see the difference in how LLM-Craft chose to change the starting state to reach these semantically different goals.
V Conclusion
@@
In this work, we present LLM-Craft, to the best of our knowledge the first LLM-based robotic shape crafting system. Through real-world experiments, we demonstrate the powerful planning capabilities of LLMs out-of-the-box for both single-step and longer horizon shape goals. However, we find that while the semantic reasoning capabilities of LLMs are very powerful, LLM-based systems sometimes hallucinate bad actions. In the LLM-Craft system there is no real-world grounding, meaning sometimes the LLM selects actions and the explanation of the predicted effect of the action does not make any physical sense. However, we believe the semantic reasoning capabilities of LLMs could be very useful in future work combining the more classical and physics-based approaches with LLMs for the higher-level reasoning and semantic similarity analysis.